7-5.324 L of a gas is collected at 25.0C and 846.0 mmHg. When the gas is changed to STP conditions, what is the new volume?
Answer #1

5) The new volume is calculated:

V2 = V1 * P1 * T2 / P2 * T1 = 5.324 L * 846 mmHg * 273 K / 760 mmHg * 298 K = 5.429 L

6) The moles of reagents are calculated:

n Glu = m / MM = 18 g / 180 g / mol = 0.1 mol

n O2 = P * V / R * T = 1 * 7.5 L / 0.082 * 310 = 0.3 mol

The limit reagent is O2, the mass of water produced is calculated:

m H2O = 0.3 mol O2 * (6 mol H2O / 6 mol O2) * (18 g H2O / 1 mol) = 5.4 g

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.